OpenStudy (anonymous):

Exponents: write with no denominaators. ax^2/b^3y

OpenStudy (anonymous):

Remember that 1/a^b=a^(-b)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now move the expressions in the denomintor up to the numerator using that

OpenStudy (anonymous):

ax^2&b^-1y?

OpenStudy (anonymous):

What happened to be being cubed? Why didn't the exponent on y not change when bringing it up the numerator?

OpenStudy (anonymous):

ug ok ax^2*b^-#y

OpenStudy (anonymous):

B^-3y,,,

OpenStudy (anonymous):

Watch that shift key. You still need to negate the exponent on y

OpenStudy (anonymous):

ax^2*b^-3-y

OpenStudy (anonymous):

Try using parenthesis to clarify. Why is y being subtracted now?

OpenStudy (anonymous):

i thought i had to negate it

OpenStudy (anonymous):

Not negate y. But the exponent of y

OpenStudy (anonymous):

Y^-1???

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Because y^(-1)=1/y, right.

OpenStudy (anonymous):

ax^2*b^-3^-3-3y-1

OpenStudy (anonymous):

i think i get it,,, maybe